The pH of a 0.15 M solution of H2CO3 is 5.52 please find the following: Show all work

  1. The first ionization expression 
  • H2CO3 + H2O ---->  H+  + HCO3-  
  1. The Ka expression
  • Ka= [H+][HCO3-]/ [H2CO3]
  1. Solve for Ka
  • 105.52 = 
  1. Solve for Kb
1
Expert's answer
2021-03-31T08:00:27-0400

2.51 × 10^24/ 6.022 × 10^23

= 0.4168 × 10^1 = 4.618mol Pb

= 4.618 × (207.2 g/mol)

= 956.8496g of lead metal